Perpetual Inventory Using FIFO

Beginning inventory, purchases, and sales for Item HM46 are as follows:

August 1 Inventory 76 units @ $20
9 Sale 64 units
13 Purchase 79 units @ $23
28 Sale 19 units

Assuming a perpetual inventory system and using the first-in, first-out (FIFO) method, determine (a) the cost of merchandise sold on August 28 and (b) the inventory on August 31.

a. Cost of merchandise sold on August 28 $__________
b. Inventory on August 31 $__________
Answer #1

a) Cost of merchandise sold on August 28 = (12*20+7*23) = $401

b) Inventory on August 31 = 72*23 = $1656